against the wind a small plane flew 21 miles in 1 hr and 10 minutes, the return trip only took 50 minutes. what was the speed of the wind, what was the speed of the plane in still air
Answer by mananth(16949)

against wind 1 1/6 hours
with wind 5/6 hours

Distance = same= 21 miles
plane speed =x
wind speed =y
t=d/r
21/(x-y)=1.17
1.17(x-y)=21.00
1.17x-1.17y=21 ....................1
21/(x+y)= 5/6
5/6(x+y )=21
5/6x+ 5/6y=21 ...............2
Multiply (1) by 5
Multiply (2) by 8
we get
5.83x-5.83y=105
6.67x+6.67y=168
12.5x=273
/12.5
x=21.84 mph plane speed

plug value of x in (1)
1.17x-1.17y=21
25.48-1.17y=21
-1.17y=21-25.48
-1.17y=-4.48
y=3.84 mph wind speed